# latexText page 4 contains only floats

### v2

@@ 3.4 Generalization Across Regions
To evaluate the generalization capability of the QuakeMetaFormer model, we divided the Turkey-L dataset into four major regions: Kahramanmaraş, Adıyaman, Osmaniye, and Malatya. The locations of the training and testing regions are illustrated in Figure 5. For this experiment, we trained two separate models on the Kahramanmaraş region (i) QuakeMetaFormer with metadata and imagery, and (ii) QuakeImageFormer with imagery only. These two trained models were then tested on the remaining three regions (see Table 6). This approach allows us to assess not only how well a model trained in one region generalizes to others, but also to determine whether adding metadata enhances generalization. This experiment will enable us to evaluate the model’s effectiveness in real earthquake scenarios, highlighting its potential for practical applications.
@@
Kah-QMF-Adı Adıyaman QuakeMetaFormer Testing
Kah-QIF-Adı Adıyaman QuakeImageFormer Testing
@@ 4.1 Impact of Adding Metadata on Performance: QuakeMetaFormer
The objective of this experiment is to assess the impact of integrating metadata with satellite imagery in improving the performance of PDA framework. The results, as shown in the Table 8, demonstrate that incorporating metadata enhances model performance across all metrics.
@@
QMF-Turkey-L 0.77 0.58 0.47 0.49 0.82
QIF-Turkey-L 0.70 0.48 0.43 0.45 0.76
QMF-Turkey-L-M 0.75 0.55 0.41 0.45 0.79
QMF-Turkey-L1 0.76 0.53 0.44 0.47 0.81
QMF-Turkey-L1-M 0.75 0.52 0.43 0.46 0.78
QMF-Turkey-L1-M-DPM 0.76 0.53 0.43 0.46 0.80
QMF-Turkey-S 0.77 0.74 0.49 0.56 0.85
QIF-Turkey-S 0.71 0.68 0.46 0.51 0.79
QMF-Turkey-S-M 0.75 0.70 0.46 0.53 0.81
QMF-Turkey-S achieved 77% accuracy compared to 71% for the QIF-Turkey-S model, while for the Turkey-L dataset, accuracy increased from 70% to 77%. In both sets of experiments, adding metadata resulted in a 7% accuracy improvement. Beyond accuracy, AUC-ROC and F1-score offer a clearer evaluation of model performance, especially in imbalanced datasets where some damage classes are underrepresented. For instance, the QMF-Turkey-S model achieved an AUC-ROC of 0.85, compared to 0.79 for the QIF-Turkey-S model, and an F1-score of 0.56, compared to 0.51 for QIF-Turkey-S. These improvements indicate that incorporating metadata significantly enhances the model’s ability to differentiate between multiple damage classes, providing a more reliable performance across all damage categories. The AUC-ROC learning curve (see Figure 6) for Turkey-L dataset shows the consistent improvement in performance for the models trained with metadata and without metadata for 150 epochs. QuakeMetaFormer reaches a 6% higher AUC-ROC than the QIF model, underscoring the importance of metadata in enhancing the model’s capacity for classifying damage more accurately. Another important observation is that adding metadata improves precision by 10% in the case of Turkey-L and 6% in the case of Turkey-S, demonstrating that incorporating metadata helps the model reduce false positives. This means the model becomes more reliable in correctly identifying buildings with actual damage, minimizing the number of undamaged buildings mistakenly classified as damaged. Notably, adding metadata not only improves the accuracy but also enhances the model’s ability to interpret imagery. For instance, QMF-Turkey-L1 achieves an AUC-ROC of 0.81, compared to QMF-Turkey-L1-M’s 0.78, reflecting only a minimal decrease in performance. Similarly, QMF-Turkey-L1-M-DPM, which simulates a scenario where only DPM data is masked during testing, achieves an AUC-ROC of 0.80. The drop in AUC-ROC is minor compared to the case of QIF-Turkey-L1, where no metadata is added during training. This experiment underscores the model’s applicability and robustness in practical scenarios where some or no metadata may be unavailable. This is particularly useful for rapid post-earthquake damage assessment, where high-latency data, such as SAR-derived parameters, may be delayed.
@@
We evaluated the effect of incorporating aftershock metadata (seismic intensity indicators) by comparing QMF-Turkey-L, which includes aftershock data, with QMF-Turkey-L1, which uses only the main earthquake event data. The inclusion of aftershock metadata resulted in a 4% increase in precision, demonstrating its role in reducing false positives. However, other metrics such as AUC-ROC and F1 showed an improvement of 1%. To understand the combined impact of multiple seismic events across different regions, we plotted the Modified Mercalli Intensity (MI) for the top 5 earthquakes and their combined MI. Figure 7 highlights that in the Turkey-Syria earthquake, the main event impacted the largest area, explaining the limited increase in model effectiveness after incorporating aftershock metadata.
@@
Figure 8: Evaluation Map for the city of Kahramanmaraş: (a) Ground Truth, (b) Predicted Values, and (c) Difference
We plotted the evaluation map for the city of Kahramanmaraş using the QMF-Turkey-S model on the test dataset to assess the extent of misclassification (see Figure 8). The evaluation map is also useful for identifying the worst-hit regions after an earthquake. The map includes three key components: the true damage class of each building, the predicted damage class, and the absolute difference between the true and predicted values. The misclassification analysis reveals that most buildings are accurately classified, with 76.7% of predictions being correct. Additionally, 17.7% of the predictions were off by ±1 class, and 1.2% by ±2 classes. Larger errors, with a difference of ±3 and ±4 classes, occurred in only 3.6% and 0.8% of cases, respectively. This analysis demonstrates the model’s robustness in predicting building-level damage with very few significant misclassifications.
@@ 4.2 Comparison of QuakeMetaFormer and Machine Learning Frameworks
Figure 9: ROC curve for QMF-Turkey-S
We evaluate the effectiveness of our proposed models with state-of-the art models published in the literature. Xiao Y et al. Yu et al. (2024) proposed a machine learning approach using a random forest classifier trained on metadata from the Turkey-S dataset. The model achieved an AUC-ROC of 0.69, outperforming DPM based methods by 11.25%. We train our proposed QuakeMetaFormer with the entire Turkey-S dataset including the satellite imagery and the results are shown in Table 9. Notably, QuakeMetaFormer achieves an AUC-ROC of 0.85, which is 16% higher than the machine learning approach (RF-Xiao). We also plot the ROC curve to understand the model’s performance across each damage class see Figure 9). The QuakeMetaFormer model demonstrated consistent performance across all classes, with a lowest AUC-ROC of 0.79 for class 3 (Needs to be demolished), indicating greater difficulty in distinguishing this class compared to others.
Similarly, Rao et al. demonstrated an ensemble approach using random forest for multiclass building damage classification across multiple datasets and reported an average F1 score of 0.29 Rao et al. (2023). In contrast, the QuakeMetaFormer models achieved an average F1 score of 0.51 across Turkey-S and Turkey-L dataset, demonstrating a significant improvement in multiclass classification. This comparison highlights that the QuakeMetaFormer architecture not only enhances overall model performance but also provides more balanced results across all damage classes, making it more reliable for building-level damage assessment.
@@ 4.3 Feature Importance Across Metadata
Using SHAP, we quantify the contribution of each metadata feature and satellite imagery to the model’s predictions. Figure 10 illustrates the impact of all input features, including imagery, on the model’s performance across four damage classes: Slightly Damaged, Heavily Damaged, Needs to be Demolished, and Collapsed. In the bar plot, higher values indicate features with a stronger influence on the model’s predictive accuracy. As anticipated, satellite imagery demonstrates the highest contribution, as its detailed representation of visible building damage plays a crucial role in decision-making, far outweighing the influence of metadata.
@@
SAR-derived parameters exert a smaller influence on predictions. Within the group, the damage proxy is more informative than SLC amplitude SAR data data, likely due to bitemporal data providing more insights into structural changes before and after the earthquake. Comparing SLC amplitude SAR data data, the results suggest that VH polarization contributes more to decision-making than VV polarization, indicating VH’s higher sensitivity to structural changes in buildings. VS30, which represents shear wave velocity, contributes little to the model; this may be due to its limited spatial variability at the building level, which reduces its effectiveness for multiclass damage classification. The limited contribution of SAR-derived parameters and VS30, compared to other features, may result from the low spatial resolution of these data sources, which fails to capture building-level changes necessary for accurate damage prediction.
@@ 4.4 Generalization Across Regions
Figure 12 summarizes the results of our experiments, aimed at evaluating the model’s ability to generalize across regions. The findings indicate that the model trained with both metadata and imagery (QMF) consistently achieved high AUC-ROC and F1 scores, demonstrating strong generalization to unseen regions. For instance, the QMF model trained on Kahramanmaraş region and tested on Adıyaman region, achieved an AUC-ROC of 0.79 and an F1 score of 0.45, while achieving 0.81 and 0.47, respectively, when both training and testing were conducted in Kahramanmaraş. This trend continues in Osmaniye and Malatya, where the QMF model demonstrates a stronger ability to generalize to unseen regions. We observe that testing the QMF model on unseen regions resulted in an average reduction of only 0.026 in AUC-ROC and 0.04 in F1 score, highlighting the model’s strong generalization ability across different regions while maintaining performance.
Figure 12: AUC-ROC (left) and F1 (right) scores across regions (y-axis) with models (QMF and QIF) on the x-axis. The QMF model demonstrates better generalization across regions and improved performance when metadata is included.
We also trained the model without metadata (QIF) to assess the impact of adding metadata on generalization. As shown in Figure 12, the QIF model experienced a significantly larger drop in performance across unseen regions, with an average reduction of 0.12 in AUC-ROC and 0.08 in F1 score, compared to QMF model (with metadata). The results show an average improvement of 15.8% in AUC-ROC and 22.8% in F1 score when metadata is included. This substantial difference highlights the positive contribution of metadata to the model’s generalization ability, making it more robust and adaptable for practical applications.
@@ 5 Conclusion
In this study, we proposed a novel framework for post-earthquake preliminary damage assessment (PDA), achieving state-of-the-art performance for the 5-class damage classification task. This problem has received limited attention in the literature to-date because of the very challenging nature of making multiclass assessments after earthquakes. Our framework achieved an AUC-ROC of 0.85, surpassing the previous leading PDA framework by 16%. Notably, the performance was consistent across all damage classes, with a minimum AUC-ROC of 0.79 for damage class 3 (Buildings that need to be demolished), and maximum AUC-ROC of 0.91 for damage class 2 (Buildings heavily damaged). We achieved this through our proposed framework, QuakeMetaFormer (QMF), which integrates high-resolution post-earthquake satellite imagery with metadata, including ground motion, SAR, and soil data. Our key contributions include (i) the development of an earthquake-specific multiclass dataset at the building level, mapped with metadata, (ii) QuakeMetaFormer, and (iii) a class-wise detailed comparative analysis of the impact of various metadata features on building-level damage assessment. To evaluate the framework, we conducted four experiments: analyzing the impact of metadata on model accuracy, comparing QuakeMetaFormer with other machine learning frameworks, assessing feature importance across metadata, and testing QuakeMetaFormer’s generalization across regions. Our results show that adding metadata to satellite imagery enhances accuracy, yielding a 7% improvement over the transformer model trained without metadata. Additionally, the model achieves an AUC-ROC of 0.8 even without high-latency metadata like SAR, demonstrating its robustness in scenarios with limited or no metadata. Compared to other ML-based methods, QuakeMetaFormer demonstrated superior performance in multiclass damage assessment, achieving an average F1-score of 0.51, significantly higher than the 0.29 achieved by competing approaches. We also conducted a detailed analysis to evaluate and understand the influence of individual metadata features on the model’s predictive performance, utilizing SHapley Additive exPlanations (SHAP). This analysis revealed that seismic indicators such as PGV, PGA, and SAR-derived parameters like DPM have a substantial impact across damage classes as compared to other parameters. We further examined how various metadata features influence model predictions at the class level, providing valuable insights into QuakeMetaFormer’s decision-making processes and its association of feature values with the extent of building damage. Lastly, incorporating metadata significantly enhanced the model’s generalization capability, improving performance when applied to new regions. The results demonstrated that adding metadata improved AUC-ROC by an average of 15.8% and F1 score by 22.8%. These improvements underscore the significant potential of our framework to accelerate disaster recovery efforts, including prioritizing inspection regions, streamlining the federal aid process, and estimating recovery costs.

### v3

@@ 3.4 Generalization Across Regions
To evaluate the generalization capability of the QuakeMetaFormer model, we divided the Türkiye-L dataset into four major regions: Kahramanmaraş, Adıyaman, Osmaniye, and Malatya. The locations of the training and testing regions are illustrated in Figure 6. For this experiment, we trained two separate models on the Kahramanmaraş region (i) QuakeMetaFormer with metadata and imagery, and (ii) QuakeImageFormer with imagery only. These two trained models were then tested on the remaining three regions (see Table 6). As discussed earlier, the similarity in data distribution between training and testing regions greatly impacts a model’s generalization capabilities, models tend to generalize better when the test data closely matches the training distribution Quiñonero-Candela et al. (2022). This experiment aims to test our hypothesis that incorporating metadata can enhance generalization, particularly by providing additional information that may help bridge distribution gaps across various regions.
@@
KAH-QMF-ADİ Adıyaman QuakeMetaFormer Testing
KAH-QIF-ADİ Adıyaman QuakeImageFormer Testing
@@ 4.1 Impact of Adding Metadata on Performance: QuakeMetaFormer
The objective of this experiment is to assess the impact of integrating metadata with satellite imagery in improving the performance of the PDA framework. The results, as shown in Table 8, demonstrate that incorporating metadata enhances model performance across all metrics.
@@
QMF-Türkiye-L 0.77 0.58 0.47 0.49 0.82
QIF-Türkiye-L 0.70 0.48 0.43 0.45 0.76
QMF-Türkiye-L-M 0.75 0.55 0.41 0.45 0.79
QMF-Türkiye-L-KAH 0.79 0.61 0.48 0.50 0.83
QMF-Türkiye-L1 0.76 0.53 0.44 0.47 0.81
QMF-Türkiye-L1-M 0.75 0.52 0.43 0.46 0.78
QMF-Türkiye-L1-M-SAR 0.76 0.53 0.43 0.46 0.80
QMF-Türkiye-S 0.77 0.74 0.49 0.56 0.85
QIF-Türkiye-S 0.71 0.68 0.46 0.51 0.79
QMF-Türkiye-S-M 0.75 0.70 0.46 0.53 0.81
QMF-Türkiye-S achieved 77% accuracy compared to 71% for the QIF-Türkiye-S model, while for the Türkiye-L dataset, accuracy increased from 70% to 77%. In both sets of experiments, adding metadata resulted in a 7% accuracy improvement. Beyond accuracy, AUC-ROC and F1-score offer a clearer evaluation of model performance, especially in imbalanced datasets where some damage classes are underrepresented. For instance, the QMF-Türkiye-S model achieved an AUC-ROC of 0.85, compared to 0.79 for the QIF-Türkiye-S model, and an F1-score of 0.56, compared to 0.51 for QIF-Türkiye-S. These improvements indicate that incorporating metadata significantly enhances the model’s ability to differentiate between multiple damage classes, providing a more reliable performance across all damage categories. The AUC-ROC learning curve (see Figure 7) for Türkiye-L dataset shows the consistent improvement in performance for the models trained with metadata and without metadata for 150 epochs. QuakeMetaFormer reaches a 6% higher AUC-ROC than the QIF model, underscoring the importance of metadata in enhancing the model’s capacity for classifying damage more accurately. Another important observation is that adding metadata improves precision by 10% in the case of Türkiye-L and 6% in the case of Türkiye-S, demonstrating that incorporating metadata helps the model reduce false positives. This means the model becomes more reliable in correctly identifying buildings with actual damage, minimizing the number of undamaged buildings mistakenly classified as damaged. Notably, adding metadata not only improves the accuracy but also enhances the model’s ability to interpret imagery. For instance, QMF-Türkiye-L1 achieves an AUC-ROC of 0.81, compared to QMF-Türkiye-L1-M’s 0.78, reflecting only a minimal decrease in performance. Similarly, QMF-Türkiye-L1-M-SAR, which simulates a scenario where SAR derived data is masked during testing, achieves an AUC-ROC of 0.80. The drop in AUC-ROC is minor compared to the case of QIF-Türkiye-L1, where no metadata is added during training. This experiment underscores the model’s applicability and robustness in practical scenarios where where high-latency data, such as SAR derived parameters (SAR-VV, SAR-VH, and DPM) may be delayed.
@@
We evaluated the effect of incorporating aftershock metadata (seismic intensity indicators) by comparing QMF-Türkiye-L, which includes aftershock data, with QMF-Türkiye-L1, which uses only the main earthquake event data. The inclusion of aftershock metadata resulted in a 5% increase in precision, demonstrating its role in reducing false positives. However, other metrics such as AUC-ROC and F1 showed an improvement of 1% and 2% respectively. Additionally, we evaluated the QMF model (QMF-Türkiye-L-KAH) on building only located in the Kahramanmaraş region to further investigate the effect of incorporating aftershock metadata for the three strongest shaking events (M 7.8 in Pazarcık, M 7.5 in Elbistan and M 6.7 Nurdağı, all located in Kahramanmaraş Province). Relative to the baseline QMF-Türkiye-L1, QMF-Türkiye-L-KAH showed an additional 2% increase in AUC-ROC and 3% in F1 score. To assess how multiple earthquakes collectively influence damage prediction across regions, we plotted the Modified Mercalli Intensity (MI) for the five largest events and their cumulative MI (Figure 8). The first three events (each >=>=> = M 6.7) produced larger overlapping region of high‐intensity aftershocks in Kahramanmaraş region, leading to a pronounced combined effect on building damage. In contrast, the subsequent aftershocks (<=<=< = M 6.7) generated minimal spatial overlap and thus contributed little additional damage comparatively. This experimental analysis explains the importance of incorporating aftershock metadata and how varying degrees of aftershock overlap affect different regions.
@@
Figure 9: Evaluation Map for the city of Kahramanmaraş (QMF-Türkiye-S): (a) Ground Truth, (b) Predicted Values, and (c) Difference
We plotted the evaluation map for the city of Kahramanmaraş using the QMF-Türkiye-S model on the test dataset to assess the extent of misclassification (see Figure 9). The evaluation map is also useful for identifying the worst-hit regions after an earthquake. The map includes three key components: the true damage class of each building, the predicted damage class, and the absolute difference between the true and predicted values. The misclassification analysis reveals that most buildings are accurately classified, with 76.7% of predictions being correct. Additionally, 17.7% of the predictions were off by ±1 class, and 1.2% by ±2 classes. Larger errors, with a difference of ±3 and ±4 classes, occurred in only 3.6% and 0.8% of cases, respectively. This analysis demonstrates the model’s robustness in predicting building-level damage with very few significant misclassifications.
@@ 4.2 Comparison of QuakeMetaFormer and Machine Learning Frameworks
Figure 10: ROC curves for the (a) QMF-Türkiye-S and (b) QMF-Türkiye-L
We evaluate the effectiveness of our proposed models with state-of-the-art models published in the literature. Yu et al. (2024) proposed a machine learning approach using a random forest classifier trained on metadata from the Türkiye-S dataset. The model achieved an AUC-ROC of 0.69, outperforming DPM based methods by 11.25%. We train our proposed QuakeMetaFormer with the entire Türkiye-S dataset including the satellite imagery and the results are shown in Table 9. Notably, QuakeMetaFormer achieves an AUC-ROC of 0.85, which is 16% higher than the machine learning approach (RF-Yu). We also plot the ROC curve to understand the model’s performance across each damage class (see Figure 10a). The QuakeMetaFormer model demonstrated consistent performance across all classes, with a lowest AUC-ROC of 0.79 for class 3 (Needs to be demolished), indicating greater difficulty in distinguishing this class compared to others. A similar trend can be observed for QMF-Türkiye-L in Figure 10b. It is important to note that a direct class-wise comparison of ROC scores between the two QuakeMetaFormer models is not appropriate, as they were trained on different datasets with varying numbers of classes, different set of buildings for training and validation, and varying input modalities. However, the results on both datasets highlights that QuakeMetaFormer architecture not only enhances overall model performance but also provides more balanced results across all damage classes, making it more reliable for building-level damage assessment in practice.
Similarly, Rao et al. (2023) evaluated their ensemble random forest framework for multiclass building damage classification across 4 earthquake events: the April 2015 Mw 7.8 Gorkha earthquake in Nepal, the September 2017 Mw 7.1 Puebla earthquake in Mexico, the January 2020 Mw 6.4 Puerto Rico earthquake, and the March 2020 Mw 5.3 Zagreb earthquake in Croatia. For multiclass classification, the reported F1 scores for these events were 0.34, 0.24, 0.23, and 0.33, respectively, with an average F1 score of 0.29. In contrast, the QuakeMetaFormer models achieved an average F1 score of 0.51 across the Türkiye-S and Türkiye-L datasets, demonstrating a significant improvement in multiclass classification (see Table 9). Although the models are not evaluated on the same event, making direct comparison challenging, the substantial difference in F1 scores (0.29 vs. 0.51) do highlights that existing machine learning methods face significant challenges in multiclass classification at the building level.
@@ 4.3 Feature Importance Across Metadata
SHAP quantifies the contribution of metadata features and satellite imagery to model predictions and confidence in class assignments. Higher SHAP importance indicates stronger influence. However, higher value doesn’t always mean significant accuracy drop when a feature is removed, as other features can compensate. Figure 11 illustrates the impact of all input features, including imagery, on the model’s performance across four damage classes: Slightly Damaged, Heavily Damaged, Needs to be Demolished, and Collapsed. In the bar plot, as anticipated, satellite imagery demonstrates the highest contribution, as its detailed representation of visible building damage plays a crucial role in decision-making, far outweighing the influence of metadata.
@@
SAR-derived parameters exert a smaller influence on predictions. Within the group, the damage proxy is more informative than SLC amplitude SAR data, likely due to bitemporal data providing more insights into structural changes before and after the earthquake. Comparing SLC amplitude SAR data, the results suggest that VH polarization contributes more to decision-making than VV polarization, indicating VH’s higher sensitivity to structural changes in buildings. VS30, which represents shear wave velocity, contributes little to the model; this may be due to its limited spatial variability at the building level, which reduces its effectiveness for multiclass damage classification. The limited contribution of SAR-derived parameters and VS30, compared to other features, may result from the low spatial resolution of these data sources, which fails to capture building-level changes necessary for accurate damage prediction.
@@ 4.4 Generalization Across Regions
Figure 13 summarizes the results of our experiments, aimed at evaluating the model’s ability to generalize across different regions affected by an earthquake event. The findings indicate that the model trained with both metadata and imagery (QMF) consistently achieved high AUC-ROC and F1 scores, demonstrating strong generalization to unseen regions. For instance, the QMF model trained on Kahramanmaraş region and tested on Adıyaman region, achieved an AUC-ROC of 0.79 and an F1 score of 0.45, while achieving 0.81 and 0.47, respectively, when both training and testing were conducted in Kahramanmaraş. This trend continues in Osmaniye and Malatya, where the QMF model demonstrates a stronger ability to generalize to unseen regions. We observe that testing the QMF model on unseen regions resulted in an average reduction (Train performance - Avg. performance on test) of only 0.027 in AUC-ROC and 0.04 in F1 score, highlighting the model’s strong generalization ability across different regions while maintaining performance.
Figure 13: AUC-ROC (left) and F1 (right) scores across regions (y-axis) with models (QMF and QIF) on the x-axis.
We also trained the model without metadata (QIF) to assess the impact of adding metadata on generalization. As shown in Figure 13, the QIF model experienced a significantly larger drop in performance across unseen regions, with an average reduction of 0.11 in AUC-ROC and 0.08 in F1 score, compared to QMF model (with metadata). The results show an average relative improvement of 15.7% in AUC-ROC and 21.8% in F1 score when metadata is included (see Table 10). This difference highlights the positive impact of metadata on the model’s generalization ability to handle differences in data distribution between training and unseen testing regions, which is not as effectively achieved when using imagery alone, making it more robust and adaptable for practical applications.
Table 10: Average performance of QMF and QIF trained on Kahramanmaraş and tested across Adıyaman, Osmaniye, and Malatya.
Model/Metric AUC-ROC F1-Score
QMF (imagery + metadata) 0.783 0.430
QIF (only imagery) 0.677 0.353
Difference (QMF - QIF) 0.106 0.077
Relative Improvement (%) 15.7% 21.8%
@@ 5 Conclusion
In this study, we proposed a novel framework for post-earthquake preliminary damage assessment (PDA), achieving state-of-the-art performance for the 5-class damage classification task. This problem has received limited attention in the literature because of the very challenging nature of making multiclass assessments after earthquakes.
Our framework achieved an AUC-ROC of 0.85, surpassing the previous leading PDA framework by 16%. Notably, the performance was consistent across all damage classes, with a minimum AUC-ROC of 0.79 for damage class 3 (Buildings that need to be demolished), and maximum AUC-ROC of 0.91 for damage class 2 (Buildings heavily damaged). We achieved this by addressing the key challenges of PDA by focusing on two critical components of the framework: diversifying data sources and enhancing data-processing approach. Instead of relying on a single modality, our proposed QuakeMetaFormer (QMF) integrates high-resolution post-earthquake satellite imagery with metadata, including ground motion, SAR, and soil data. To evaluate the framework, we conducted four experiments: analyzing the impact of metadata on model accuracy, comparing QuakeMetaFormer with other machine learning frameworks, assessing feature importance across metadata, and testing QuakeMetaFormer’s generalization across regions.
Our results show that adding metadata to satellite imagery enhances accuracy, yielding a 7% improvement over the transformer model trained without metadata. Additionally, the model achieves an AUC-ROC of 0.8 even without high-latency metadata like SAR, demonstrating its robustness in scenarios with limited or no metadata.
Compared to other ML-based methods, QuakeMetaFormer demonstrated superior performance in multiclass damage assessment, achieving an average F1-score of 0.51, significantly higher than the 0.29 achieved by competing approaches. We also conducted a detailed analysis to evaluate and understand the influence of individual metadata features on the model’s predictive performance, utilizing SHapley Additive exPlanations (SHAP). This analysis revealed that seismic indicators such as PGV, PGA, and SAR-derived parameters like DPM have a substantial impact across damage classes as compared to other parameters.
Notably, incorporating metadata significantly enhanced the model’s generalization capability, improving performance across different regions affected by the same earthquake event. The results demonstrated that adding metadata improved AUC-ROC by an average of 15.7% and F1 score by 21.8%.
These improvements underscore the significant potential of our framework to accelerate disaster recovery efforts, including prioritizing inspection regions, streamlining the federal aid process, and estimating recovery costs.
6 Future Work
While the proposed PDA framework demonstrates state-of-the-art performance and generalization across different regions affected by an earthquake, two important challenges remain open for future research. First, generalization across distinct earthquake events, with varying building typologies, regional conditions, and seismic characteristic, has not yet been fully explored. Future work should validate the framework on multiple disaster datasets to assess robustness under diverse real-world scenarios. Lastly, the potential of low latency high-resolution SAR imagery should be further explored to enhance the accuracy and effectiveness of PDA frameworks.